OpenStudy (anonymous):

Can Some one PLEASE help me wit this!!!!!?!?!??? Graph the linear function by plotting the x- and y-intercepts. 5x - 25y - 25 = 0

OpenStudy (anonymous):

x and y can be anything

OpenStudy (anonymous):

how would i find the x and y intercepts

OpenStudy (anonymous):

it can be 4, 5, 6, 14, 2? idk sry what do i know

OpenStudy (anonymous):

Thx anyway. U really helped.

OpenStudy (anonymous):

um np, for no problem sry i couldn't help that much

OpenStudy (anonymous):

5x - 2y - 25 = 0 5(0) - 2y - 25 = 0 -2y - 25 = 0 -2y = 25 y = -25/2 < -- y intercept 5x - 2y - 25 = 0 5x - 2(0) - 25 = 0 5x - 0 - 25 = 0 5x - 25 = 0 5x = 25 x = 5 < -- x intercept

OpenStudy (anonymous):

Thank you. You are a lifesaver.